OpenStudy (anonymous):

evalute the integral (sin^3 (x)/cosx)dx

OpenStudy (jamesj):

Note that sin^3 = sin(1- cos^2) Put that into your expression and it should look quite a bit easier.

OpenStudy (zarkon):

\[\frac{\sin^3(x)}{\cos(x)}=\frac{\sin(x)\sin^2(x)}{\cos(x)}=\frac{\sin(x)(1-\cos^2(x))}{\cos(x)}=\cdots\]

OpenStudy (anonymous):

can u continue plz, because i needed to check the answer

OpenStudy (anonymous):

and is integral

OpenStudy (jamesj):

Tell me what you think the answer is and I'll tell you if it's right. First the integral of sin x/cos x. And then the integral of sin x.cos^2 x/cos x

OpenStudy (anonymous):

you cant do that i guess... do we put u= cosx ?

OpenStudy (jamesj):

In the first one, yes.

OpenStudy (anonymous):

and then is the - integral (1-u^2)/u du ??

OpenStudy (jamesj):

No. if u = cos x, then du/dx = -sin x so \[\frac{\sin x}{\cos x} dx = - \frac{du/dx}{u} dx = - \frac{du}{u}\]

OpenStudy (anonymous):

Is the answer >

OpenStudy (anonymous):

-ln|cosx| + (cos^2(x)/2) ?

OpenStudy (jamesj):

And make sure you add a constant C.

OpenStudy (anonymous):

Isnt the sign positive.. becuz the - would multiply in integral

OpenStudy (jamesj):

yes, you're right.

OpenStudy (anonymous):

put u=cosx du=-sinxdx so -(1-u^2)/u du -(ln(u)-u^2/2)+c ln(sec(x))+cosx^2/2+c